OpenStudy (anonymous):

A rock dropped from an unknown height hits the ground below with a velocity of 15.0 m/s. From what height was the rock dropped?

OpenStudy (jiteshmeghwal9):



OpenStudy (jiteshmeghwal9):

the question is a case of motion under gravity. step 1 is taking sign convention . I take upward direction positive & downward direction negative. so, a=-g u=initial velocity=0 v=final velocity=15 m/s h=height from which the ball was dropped.

OpenStudy (jiteshmeghwal9):

use the equation \(v^2=u^2-2gh\)

OpenStudy (jiteshmeghwal9):

i have given u all the values. substitute the values in the equation & gt ur answer .

OpenStudy (anonymous):

Thank you!

OpenStudy (jiteshmeghwal9):

yw :)